OpenStudy (anonymous):

Which of the following describes the correct process for solving the equation 2x - 6 = 22 and arrives at the correct solution? A. Divide both sides by -6 and then add 2. The solution is x = 8. B. Subtract 6 from both sides of the equation and then divide by 2. The solution is x = 7. C. Add 6 to both sides of the equation and then divide by 2. The solution is x = 14. D. Add 6 to both sides of the equation and then divide by 22. The solution is x = -14.

OpenStudy (anonymous):

@Jacob902 i think its either C or D

Mehek (mehek14):

ok so the first part of C and D is correct If you did that, your equation would be \(2x=28\) now the step after this would be to divide the number attached to the variable, which is 2 in your case.

OpenStudy (jacob902):

c

OpenStudy (anonymous):

ok thnx guys :)

OpenStudy (jacob902):

2x-6=22 Well first you're going to add six to both sides: 2x-6+6=22+6 2x=28 Then, you're going to divide both sides by 2: 2/2x=28/2 x=14 Select the answer you think is correct. :o)

OpenStudy (jacob902):

im postive its c

OpenStudy (anonymous):

thnx :)